OpenStudy (anonymous):

A painter used 3/8 gallon of paint for one job and 1/3 gallon for another job. How many gallons of paint did the painter use for both jobs? A. gal B. gal C. gal D. 1 gal

OpenStudy (anonymous):

A = 4/11 B = 17/24 C = 3/4 D = 1 1/8

OpenStudy (anonymous):

D= 1 1/8

OpenStudy (anonymous):

medal to one whos right

OpenStudy (anonymous):

@seb.cal

OpenStudy (anonymous):

errrr never mind its 7/24